A 600-room hotel has 476 occupied on 10/31/18. The hotel made $134,567.84.   What is the ROP, ADR and REVPAR?
Answer #1

ADR also knowns average daily rate

  • REVPAR also known as Revenue Per Available Room calculated as Room Revenue/Rooms Available = 134,537.84/600 = $ 224.22
  • ADR also known as average daily revenue is calculated as Room revenue / room sold = 134,537.84/476 = $ 282.64
  • ROP also known as rate of occupancy can be calculated by room occupied / room available = 476/600 = 79.33%